OpenStudy (anonymous):

A water tank can be filled by two pipes together in 6 minutes. if the tank is filled by the pipes individually, it would take the smaller pipe 5 minutes longer than the big pipe. Find the time in which each pipe alone would fill the tank.

OpenStudy (anonymous):

\[\left( \frac{ 1 }{ s }+\frac{ 1 }{ l } \right)\cdot 6=1\] and \[s+5 =l\] you can use substitution to solve

OpenStudy (anonymous):

do you understand how i came up with those?

OpenStudy (anonymous):

oops,\[s=l+5\]

OpenStudy (anonymous):

s is the time it takes the smaller pipe to fill the tank alone l is the time it takes to fill the tank alone 6/s is the portion of the tank filled by the smaller pipe in 6 minutes 6/l is the portion of the tank filled by the larger pipe in 6 minutes 6/s + 6/l = 1 which is the number of tanks filled in 6 minutes by both pipes

OpenStudy (anonymous):

@pgpilot326 Thank you for your help. Can you please tell me what value you have calculated for the big and small pipe?

OpenStudy (anonymous):

i haven't calculated them, have you?